Magical Mathematics[Interesting Approach]

a1,a2,a3,.....,an,....
in the sequence above,each term after the first term is equal to the preceding term plus the constant c,if a1+a3+a5=27 what is the value of a2+a4=?

srinivas
12 Years agoGrade

1 Answer

Jitender Singh
12 Years ago
Ans: 18
Sol:
Let 1stterm a1 be a.
Series would be like this a, a + c, a + 2c, a + 3c, a + 4c, ….......
a1 + a2+ a3=a + (a + 2c) + (a + 4c) = 27
3a + 6c = 27
a + c = 9
a2+ a4= (a + c) + (a + 3c) = 2a + 4c = 2(a + 2c) = 18
Cheers!
Thanks & Regards
Jitender Singh
IIT Delhi
askIITians Faculty